# Lower Bounds for Book Ramsey Numbers

### v1

@@ 2 Results
The book Ramsey number $R(B_{n-1},B_{n})$ equals $14n-$ 1 for $n\leq 20$. Moreover, if $12n-$ 1 is a prime power congruent to $1$ modulo $4$, then $R(B_{n-1},B_{n})=4n-1$.
We also computed the following new bounds for other values $R(B_{r},B_{s})$. We note that some of these bounds, namely $R(B_{2},B_{8})\geq 21,R(B_{2},B_{9})\geq 22,$ and $R(B_{5},B_{7})\geq 25$, have been computed independently using tabu search methods in the recent paper by Lidický, McKinley, and Pfender [23]. We were also made aware that many of the same bounds given below have also been computed independently by Van Overberghe [29]. The bound $R(B_{3},B_{6})\geq 19$ also appeared in the author’s Ph.D. thesis [30].
@@
A Ramsey $(G,H,n)$ graph is a graph on $n$ vertices that does not contain a copy of $G$ or $\overline{H}$. We say that such a graph is a critical graph if $n=R(G,H)-1$. It is of general interest to enumerate the number of critical graphs (up to isomorphism) for given $G$ and $H$. The following theorem gives the numbers of critical graphs for $G=B_{r}$, $H=B_{s}$ and small values of $r$ and $s$.
@@
$r$ $s$ $R(B_{r},B_{s})$critical graphs
$3$ $4$ $15$ $1$
$4$ $5$ $19$ $27$
$5$ $6$ $23$ $23$
$2$ $8$ $21$ $1$
$2$ $9$ $22$ $72$
$2$ $10$ $25$ $5$
$3$ $6$ $19$ $4$
@@ 3 Block-circulant Graphs
For additive groups $X,Y$, let
$\Delta(X,Y,d)$ :=|$:=|\{(x,y)\in X\times Y:x-y=d\}|,$
$\Sigma(X,Y,d)$ :=|$:=|\{(x,y)\in X\times Y:x+y=d\}|.$
@@
A useful property of 2-block-circulant graphs is that $\Gamma(u,v)$ can be computed easily via the following lemma.
@@
More generally, we can make the following construction: Let $G$ be an abelian group. Let $D_{11},D_{12},D_{22}\subseteq G$. Define a graph $\Gamma_{G}(D_{11},D_{12},D_{22})$ with vertex set $V_{1}\sqcup V_{2}$ with $V_{1}=V_{2}=G$ such that $x,y$ are adjacent if and only if $x,y\in V_{1}$ and $y-x\in D_{11}$, $x,y\in V_{2}$ and $y-x\in D_{22}$, or $x\in V_{1},y\in V_{2}$ and $y-x\in D_{12}$. A 2-block-circulant graph on $2m$ vertices is the special case $G=\mathbb{Z}_{m}$.
@@
Recall that in the Paley graph of order $q$, vertices $x$ and $y$ are adjacent if and only if $x-y$ is a nonzero quadratic residue in $\mathbb{F}_{q}$. In particular, for prime $q$, Paley graphs are circulant. In Theorem 1, the lower bound $R(B_{n},B_{n})\geq 4n+2$ when $q=4n+1$ is a prime power comes from the fact that the Paley graph of order $q$ does not contain copies of $B_{n}$ and is self-complementary. More precisely, two adjacent vertices share exactly $n-1$ common neighbors. Our construction in the proof of Theorem 2 is a generalization of the Paley graphs using two blocks instead of one. The following lemma gives convenient properties of quadratic residues and nonresidues in $\mathbb{F}_{q}$.
@@
Let $q=2n-1$ be a prime power equivalent to 1 modulo 4. Let $Q$ and $N$ denote respectively the set of quadratic residues and nonresidues in $\mathbb{F}_{q}$. Set $D_{11}=D_{12}=Q$, and $D_{22}=N$. We will show that the conditions in Lemma 9 hold for the graph $\Gamma_{\mathbb{F}_{q}}(Q,Q,N)$.
Since $q\equiv 1\pmod{4}$, we have $-1\in Q$. It follows that $d\in Q$ if and only if $-d\in Q$, and also $d\in N$ if and only if $-d\in N$. Therefore $\overline{D}_{22}=D_{11}=Q$, $\overline{D}_{11}=D_{22}=N$, and $\overline{D}_{12}=N\cup\{0\}$.
@@
=$=\begin{cases}\frac{q-1}{2}+0\text{ if }d=0\\ \frac{q-1}{4}+\frac{q-1}{4}\text{ if }d\neq 0\\ \end{cases}$
@@ 4.1 Proofs of Theorem 3 and Theorem 4
The bounds in Theorem 3 and the bounds for $n\leq 20$ in Theorem 2 were computed using the SAT solver Kissat [4] and the IP solver SCIP [15]. Generally speaking, the lower bounds in Table 1 were produced using SAT solving and those in Tables 2 and 3 were produced by integer programming. For each lower bound, we produced a graph avoiding copies of $B_{r}$ and $\overline{B}_{s}$. These graphs can be found in the Appendix, and we display two such graphs in Figure 2.
@@
For the new upper bounds $R(2,8)\leq 21$, $R(2,9)\leq 22$, and $R(3,7)\leq 20$, we used Encoding B and showed the corresponding formula was unsatisfiable. For each bound we produced DRAT certificates, sequences of clauses (and deletions) that are redundant with respect to the formula. These certificates are then used to verify the computation by the SAT solver is correct. The most difficult was by far $R(3,7)\leq 20$, which yielded a DRAT certificate of size approximately 19GB and was verified in drat-trim [31] in 39292 seconds.
$r$ $s$ $R(r,s)$ Upper bound solve time (seconds)
@@
$r$ $s$ $R(B_{r},B_{s})$critical Ramsey graphs Computation time (seconds)
$3$ $4$ $15$ $1$ $1$
$4$ $5$ $19$ $27$ $931$
$5$ $6$ $23$ $23$ $985401$
$2$ $8$ $21$ $1$ $5$
$2$ $9$ $22$ $72$ $154$
$2$ $10$ $25$ $5$ $22$
$3$ $6$ $19$ $4$ $60$

### v2

@@ 2 Results
The book Ramsey number $R(B_{n-1},B_{n})$ equals 4n14n-1 for $n\leq 20$. Moreover, if 2n12n-1 is a prime power congruent to $1$ modulo $4$, then $R(B_{n-1},B_{n})=4n-1$.
We also computed the following new bounds for other values $R(B_{r},B_{s})$. We note that some of these bounds, namely $R(B_{2},B_{8})\geq 21,R(B_{2},B_{9})\geq 22,$ and $R(B_{5},B_{7})\geq 25$, have been computed independently using tabu search methods in the recent paper by Lidický, McKinley, Pfender, and Van Overberghe [23]. The bound $R(B_{3},B_{6})\geq 19$ also appeared in the author’s Ph.D. thesis [30].
@@ 99 $11$ $\bf{41}$
A Ramsey $(G,H,n)$ graph is a graph on $n$ vertices that does not contain a copy of $G$ or $\overline{H}$. We say that such a graph is a critical graph if n=R(G,H)1n=R(G,H)-1. It is of general interest to enumerate the critical graphs (up to isomorphism) for given $G$ and $H$, and this was done for $(G,H)\in\{(B_{2},B_{3}),(B_{3},B_{4}),(B_{2},B_{6}),(B_{2},B_{7})\}$ in [28, 5]. The following theorem gives the numbers of critical graphs for $G=B_{r}$, $H=B_{s}$ and small values of $r$ and $s$.
@@
$r$ $s$ $R(B_{r},B_{s})$ #critical graphs
11 11 66 11
11 22 77 44
11 33 99 88
11 44 $11$ 77
11 55 $13$ 88
11 66 $15$ 88
11 77 $17$ $10$
11 88 1919 $10$
22 22 $10$ 11
22 33 $11$ 44 [28]
22 44 $13$ 66
22 55 $16$ 11
22 66 $17$ 33 [5]
22 77 $18$ $65$ [5]
22 88 2121 11
22 99 $22$ $72$
22 $10$ $25$ 55
22 $11$ $28$ 11
22 $12$ $28$ $10$
$r$ $s$ $R(B_{r},B_{s})$ #critical graphs
33 33 $14$ 11
33 44 $15$ 11 [28]
33 55 $17$ $10$
33 66 1919 44
44 44 $17$ 11
44 55 1919 $27$
55 55 2121 $247$
55 66 $23$ $23$
66 66 $26$ $15$
@@ 3 Block-circulant Graphs
For subsets $X,Y$ of an additive group $G$ and $d\in G$, let
$\Delta(X,Y,d)$ $:=|\{(x,y)\in X\times Y:x-y=d\}|,$
$\Sigma(X,Y,d)$ $:=|\{(x,y)\in X\times Y:x+y=d\}|.$
@@
A useful property of 2-block-circulant graphs is that $|\Gamma(u,v)|$ can be computed easily via the following lemma.
@@
More generally, we can make the following construction, which is a “2-block” Cayley graph. Let $G$ be an abelian group, and let $D_{11},D_{12},D_{22}\subseteq G$. Define a graph $\Gamma_{G}(D_{11},D_{12},D_{22})$ with vertex set $V_{1}\sqcup V_{2}$ with $V_{1}=V_{2}=G$ such that $x,y$ are adjacent if and only if $x,y\in V_{1}$ and $y-x\in D_{11}$, $x,y\in V_{2}$ and $y-x\in D_{22}$, or $x\in V_{1},y\in V_{2}$ and $y-x\in D_{12}$. A 2-block-circulant graph on $2m$ vertices is the special case $G=\mathbb{Z}_{m}$.
@@
Recall that in the Paley graph of order $q$, vertices $x$ and $y$ are adjacent if and only if xyx-y is a nonzero quadratic residue in $\mathbb{F}_{q}$. In particular, for prime $q$, Paley graphs are circulant. In Theorem 1, the lower bound $R(B_{n},B_{n})\geq 4n+2$ when $q=4n+1$ is a prime power comes from the fact that the Paley graph of order $q$ does not contain copies of $B_{n}$ and is self-complementary. More precisely, two adjacent vertices share exactly n1n-1 common neighbors. Our construction in the proof of Theorem 2 is a generalization of the Paley graphs using two blocks instead of one. The following lemma gives convenient properties of quadratic residues and nonresidues in $\mathbb{F}_{q}$.
@@
Let q=2n1q=2n-1 be a prime power equivalent to 1 modulo 4. Let $Q$ and $N$ denote respectively the set of quadratic residues and nonresidues in $\mathbb{F}_{q}$. Set $D_{11}=D_{12}=Q$, and $D_{22}=N$. We will show that the conditions in Lemma 9 hold for the graph $\Gamma_{\mathbb{F}_{q}}(Q,Q,N)$.
Since $q\equiv 1\pmod{4}$, we have $1\in Q$. It follows that $d\in Q$ if and only if $d\in Q$, and also $d\in N$ if and only if $d\in N$. Therefore $\overline{D}_{22}=D_{11}=Q$, $\overline{D}_{11}=D_{22}=N$, and $\overline{D}_{12}=N\cup\{0\}$.
@@
[…]
@@ 4.1 Proofs of Theorem 3 and Theorem 4
The bounds in Theorem 3 and the bounds for $n\leq 20$ in Theorem 2 were computed using the SAT solver Kissat [4] and the IP solver SCIP [15]. Generally speaking, the lower bounds in Table 1 were produced using SAT solving and those in Tables 2 and 3 were produced by integer programming. However, there are two exceptions: the lower bounds $R(B_{2},B_{12})\geq 28$ and $R(B_{2},B_{13})$ follow from results in [14]. Namely, the Schläfli graph is a 27-vertex graph that contains no copy of $B_{11}$ and its complement contains no $B_{2}$, so $28\leq R(B_{2},B_{11})\leq R(B_{2},B_{12})$. Moreover, the complete bipartite graph $K_{14,14}$ contains no $B_{2}$ and its complement contains no $B_{13}$, so $R(B_{2},B_{13})\geq 29$. For each lower bound, we produced a graph avoiding copies of $B_{r}$ and $\overline{B}_{s}$. These graphs can be found in the Appendix, and we display two such graphs in Figure 2.
@@
For the new upper bounds $R(B_{2},B_{8})\leq 21$, $R(B_{2},B_{9})\leq 22$, and $R(B_{3},B_{7})\leq 20$, we used Encoding B and showed the corresponding formula was unsatisfiable. For each bound we produced DRAT certificates, sequences of clauses (and deletions) that are redundant with respect to the formula. These certificates are then used to verify the computation by the SAT solver is correct. The most difficult were by far $R(B_{3},B_{7})\leq 20$ and $R(B_{2},B_{13})\leq 29$, which yielded DRAT certificates of sizes approximately 19GB and 186GB, respectively.
$r$ $s$ $R(B_{r},B_{s})$ Upper bound solve time (seconds)
@@
$r$ $s$ $R(B_{r},B_{s})$ #critical Ramsey graphs Computation time (seconds)
11 11 66 11 11
11 22 77 44 11
11 33 99 88 11
11 44 $11$ 77 11
11 55 $13$ 88 11
11 66 $15$ 88 $23$
11 77 $17$ $10$ $1367$
11 88 1919 $10$ $125564$
22 22 $10$ 11 11
22 33 $11$ 44 [28] 11
22 44 $13$ 66 11
22 55 $16$ 11 11
22 66 $17$ 33 [5] 11
22 77 $18$ $65$ [5] 44
22 88 2121 11 55
22 99 $22$ $72$ $154$
22 $10$ $25$ 55 $22$
22 $11$ $28$ 11 22
22 $12$ $28$ $10$ $985$
33 33 $14$ 11 11
33 44 $15$ 11 [28] 11
33 55 $17$ $10$ 22
33 66 1919 44 $60$
44 44 $17$ 11 11
44 55 1919 $27$ $931$
55 55 2121 $247$ $4250$
55 66 $23$ $23$ $985401$
66 66 $26$ $15$ $28264$